# MOWA: Multiple-in-One Image Warping Model

### v2

@@ IV-E Generalization Evaluation
We show the generalization ability of the proposed method in terms of the cross-domain and zero-shot evaluations in Fig. 7. For the cross-domain evaluation, the new inputs belong to the above six practical image warping tasks but they are captured in real-world settings with different cameras, resolutions (can be up to 4K), and scenes. For the zero-shot evaluation, we consider a new image warping task, i.e., image retargeting, which aims to flexibly change the image scale without distorting the content as much as possible.
@@
The visualization results demonstrate that MOWA can well extend to real-world scenarios, though its training datasets are mostly synthesized by hand-crafted camera models or manipulation spaces. One possible reason is the multiple-in-one model can naturally address the overfitting issue on specific datasets by learning various tasks. In addition to the cross-domain evaluations, we find that while our model does not involve the image retargeting task during training, it is still able to warp the image based on the “content-aware” principle. As we can observe, the predicted control points are accurately aligned to the image boundaries. Such knowledge transferring to new tasks potentially benefits from the shared motion perception across different tasks. Therefore, our results show fewer geometrical distortions for the foreground than the crop and resize operation, with the face and body experiencing less stretching.
It is noticed that our method also exhibits satisfactory robustness to noisy data. For instance, in Figure 8, although MOWA is trained on the synthetic dataset with clean masks, the warping results of the real-world dataset with noisy pseudo masks (the clean masks are not available in some practical applications) are still structurally reasonable.
IV-F Limitation Discussion
@@
We show some failure cases in Figure 9. In these cases, we can find that the image boundaries are more irregular and the expected displacements of warping are more complicated than most samples. Consequently, it is challenging to approximate the accurate motion structure with a certain number of control points. This limitation could be addressed by adding more control points and cascading more TPS regression heads. Besides, scaling up the resolution of the input image could potentially improve the warping performance on image boundaries and details.
@@ V Conclusion
We have proposed MOWA in this work, the first multiple-in-one image warping framework in the field of computational photography. It considers six representative and practical tasks in one learning model and uses a unified motion representation to achieve various warping purposes. In particular, to mitigate the difficulty of approximating diverse motions of different tasks, we propose to disentangle the motion estimation at both the region level and pixel level. Then, we enable MOWA the explicit task-aware ability by introducing a lightweight point-based classifier. Compared to the common image-based classifier, it can achieve comparable performance while offering significant parameter reduction. Subsequently, we feed the task label predicted by the task classifier into a prompt learning module and further modulate the feature maps in the decoder, which facilitates a single model to efficiently navigate and leverage its extensive parameter space to meet various warping requirements. Comprehensive experiments demonstrate that MOWA outperforms different SotA methods specifically designed for each single task, with an affordable model size. In the future, we plan to empower MOWA with cross-view and cross-modal abilities, aiming to build a foundation model for universal image warping.

### v3

@@ IV-E Generalization Evaluation
We show the generalization ability of the proposed method in terms of the cross-domain and zero-shot evaluations in Fig. 7. For the cross-domain evaluation, the new inputs belong to the above six practical image warping tasks but they are captured in real-world settings with different cameras, resolutions (from 256$\times$256 to 4K), and scenes. For the zero-shot evaluation, we consider a new image warping task, i.e., image retargeting, which aims to flexibly change the image scale without distorting the content as much as possible.
@@
The visualization results demonstrate that MOWA can well extend to real-world scenarios, though its training datasets are mostly synthesized by hand-crafted camera models or manipulation spaces. One possible reason is the multiple-in-one model can naturally address the overfitting issue on specific datasets by learning various tasks. In addition to the cross-domain evaluations, we find that while our model does not involve the image retargeting task during training, it is still able to warp the image based on the “content-aware” principle. The content-awareness of the proposed warping model is learned from the training datasets in a data-driven manner, in which the labels are constructed by optimizing an energy function with line-preserving mesh deformation or gradient magnitude [24, 5, 52]. As we can observe, the predicted control points are accurately aligned to the image boundaries. Such knowledge transferring to new tasks potentially benefits from the shared motion perception across different tasks. Therefore, our results show fewer geometrical distortions for the foreground than the crop and resize operation, with the face and body experiencing less stretching. It is noticed that our method also exhibits satisfactory robustness to noisy data. For instance, in Figure 8, although MOWA is trained on the synthetic dataset with clean masks, the warping results of the real-world dataset with noisy pseudo masks (the clean masks are not available in some practical applications) are still structurally reasonable.
IV-F Extended Applications
While the proposed method is originally designed for single-image warping tasks, it can be extended to multi-view image warping applications. In particular, we add a feature matching module to the framework, e.g., cost volume, and then fine-tune the pre-trained MOWA on the multi-view warping tasks such as image alignment [53, 54], image stitching [55], and stereo image rectification [56]. The visual results of MOWA are illustrated in Figure 9. As we can observe, our method is capable of addressing image alignment and image stitching problems under different viewpoints, contrasts, and illuminations. Please note that the linearly average fusion algorithm is applied for the stitching. Better fusion results can be achieved by using other advanced composition strategies like seam cutting, but the scope of this work primarily focuses on the performance of image warping. Additionally, the proposed method can rectify the rotations of uncalibrated stereo images under different scenes. The stereo matching results (inferred by HITNet [51]) of our rectified images also show fewer artifacts and more reasonable distributions than those of uncalibrated inputs.
IV-G Limitation Discussion
@@
We show some failure cases in Figure 10 and find that the image boundaries are more irregular and the expected displacements of warping are more complicated than most samples. Consequently, it is challenging to approximate the accurate motion structure with a certain number of control points. This limitation could be addressed by adding more control points and cascading more TPS regression heads. Scaling up the resolution of the input image could also potentially improve the warping performance of image boundaries and details.
@@ V Conclusion
We have proposed MOWA in this work, the first multiple-in-one image warping framework in the field of computational photography. It considers six representative and practical tasks in one learning model and uses a unified motion representation to achieve various warping purposes. In particular, to mitigate the difficulty of approximating diverse motions of different tasks, we propose to disentangle the motion estimation at both the region level and pixel level. Then, we enable MOWA’s explicit task-aware ability by introducing a lightweight point-based classifier. Compared to the common image-based classifier, it can achieve comparable performance while offering significant parameter reduction. Subsequently, we feed the task label predicted by the task classifier into a prompt learning module and further modulate the feature maps in the decoder, which facilitates a single model to efficiently navigate and leverage its extensive parameter space to meet various warping requirements. Comprehensive experiments demonstrate that MOWA outperforms different SotA methods specifically designed for each single task, with an affordable model size. In the future, we plan to empower MOWA with cross-view and cross-modal abilities, aiming to build a foundation model for universal image warping.